# Integral canonical models of exceptional Shimura varieties

### v1

@@ 4.3 Making period maps proper
Let $F$ be a mixed charactersitic local field with ring of integers $\mathcal{O}$ and residue field $k$. Let $\mathcal{P}/\mathcal{O}$ be smooth with a partial log smooth compactification $(\mathcal{P}^{\prime},\mathcal{D})/\mathcal{O}$. Let $f:\mathcal{Z}\to\mathcal{P}$ be a smooth projective morphism with geometrically connected fibers of relative dimension $<\textrm{char}(k)$. Let $\mathbf{M}=H^{k}_{DR}(\mathcal{Z}/\mathcal{P})$ be the $p$-torsion Fontaine–Laffaille module with underlying filtered flat vector bundle the De Rham cohomology of the special fiber $(M,\nabla,F^{\bullet}M)=H^{k}_{DR}(\mathcal{Z}_{0}/\mathcal{P}_{0})$. Assume:
@@
    There is a log smooth curve (possibly not irreducible) $(\mathcal{C}^{\prime},\Delta)/\mathcal{O}$ and a morphism $g^{\prime}:(\mathcal{C}^{\prime},\Delta)\to(\mathcal{P}^{\prime},\mathcal{D})$ with $\Delta=(\mathcal{C}^{\prime}\times_{\mathcal{P}^{\prime}}\mathcal{D})^{\mathrm{red}}$ such that $g^{\prime}(\Delta)$ meets every component of $\mathcal{D}$ and the base-change $\mathcal{Z}_{\mathcal{C}}$ (where $\mathcal{C}=\mathcal{C}^{\prime}\setminus\Delta$) has a log smooth compactification $(\mathcal{Z}^{\prime}_{\mathcal{C}},\mathcal{E})/\mathcal{O}$ and a morphism $f^{\prime}:\mathcal{Z}^{\prime}_{\mathcal{C}}\to\mathcal{C}^{\prime}$ extending the natural one.
@@
    $g^{\prime*}(M^{\prime},\nabla^{\prime},F^{\bullet}M^{\prime})\cong H^{k}_{DR}((\mathcal{Z}^{\prime}_{\mathcal{C},0},\mathcal{E}_{0})/(\mathcal{C}_{0}^{\prime},\Delta_{0}))$ via an isomorphism extending the natural one, where $H^{k}_{DR}((\mathcal{Z}^{\prime}_{\mathcal{C},0},\mathcal{E}_{0})/(\mathcal{C}_{0}^{\prime},\Delta_{0}))$ is the relative log De Rham cohomology.
@@
The main step is the following straightforward claim. Throughout by “$\tilde{\mathcal{C}}$” we mean the base-change to $W_{2}(k)$.
Claim.
There is an isomorphism
$F^{*}\operatorname{gr}_{F}\mathbb{H}^{k}(\Omega^{\bullet}_{\mathcal{Z}^{\prime}_{\mathcal{C},0}/\mathcal{C}_{0}^{\prime}}(\log\mathcal{E}_{0}))\xrightarrow{\Phi}\mathbb{H}^{k}(\Omega^{\bullet}_{\mathcal{Z}_{\mathcal{C},0}/\mathcal{C}_{0}^{\prime}}(\log\mathcal{E}_{0}))$
extending the usual Frobenius morphism of Deligne–Illusie over $\mathcal{C}_{0}$.
Proof.
By log smoothness, both $\tilde{\mathcal{C}}^{\prime}$ and $\tilde{\mathcal{Z}}^{\prime}_{\mathcal{C}}$tale locally [^4] In fact, the lifts of Frobenius preserving the boundary exist Zariski locally. To simplify notation, if $(\tilde{X},\tilde{D})/W_{2}$ is log smooth and $(X,D)$ the reduction mod $p$, then there is a natural exact sequence $0\to F^{*}T_{X}(-\log D)\to F^{*}T_{X}\to\bigoplus_{i}F^{*}\mathcal{O}_{D_{i}}(D_{i})\to 0$ where the $D_{i}$ are the components of $D$ with defining equations $f_{i}$, and the last map sends $1\otimes\delta$ to $1\otimes\frac{\delta f_{i}}{f_{i}}\mod f_{i}$ on the $i$th factor. Lifts of Frobenius to $\tilde{X}$ are a torsor for $H^{0}(F^{*}T_{X})$, so if $\tilde{F}$ is a local lift of Frobenius to $\tilde{X}$, and $\tilde{f}_{i}$ local defining equations for $\tilde{D}_{i}$, then $\tilde{F}(\tilde{f}_{i})=\tilde{f}^{p}_{i}+pg_{i}$, and Zariski locally we may modify $\tilde{F}$ so that each $\tilde{F}(\tilde{f}_{i})$ is divisible by $\tilde{f}_{i}$. admit compatible lifts of Frobenius fixing each component of the boundary. Thus, as in [deligneillusie], we obtain a morphism of filtered complexes
$(\Omega_{\mathcal{Z}^{\prime}_{\mathcal{C},0}/\mathcal{C}_{0}^{\prime}}^{\bullet}(\log\mathcal{E}_{0}),0)\to(F_{*}\Omega_{\mathcal{Z}^{\prime}_{\mathcal{C},0}/\mathcal{C}_{0}^{\prime}}^{\bullet}(\log\mathcal{E}_{0}),F_{*}d)$
which restricts to the usual quasi-isomorphism over $\mathcal{C}_{0}$ and is easily checked to be a quasi-isomorphism. ∎
To prove the proposition, we must show the Frobenius $\Phi:C^{-1}_{\mathcal{P}}\operatorname{gr}_{F}\mathbf{M}\to\mathbf{M}$ extends to an isomorphism $\Phi^{\prime}:C^{-1}_{\mathcal{P}^{\prime}}(\operatorname{gr}_{F}M^{\prime},\operatorname{gr}_{F}\nabla^{\prime})\to(M^{\prime},\nabla^{\prime})$. Since both sides are locally free, it suffices to show that $\Phi$ extends in codimension one, which may therefore be checked on the pullback $g^{\prime*}(M^{\prime},\nabla^{\prime},F^{\bullet}M^{\prime})$. This is the claim. ∎
@@ 5 Existence of integral canonical models
Let $V$ denote a faithful rational representation of $G$, and $\mathbb{V}\subset V$ denote a lattice. The choice of $\mathbb{V}$ specifies a compact open subgroup of $G(\mathbb{A}_{f})$ by considering the stabilizers of $\mathbb{V}_{\hat{\mathbb{Z}}}$. We choose the level structure $K$ to be a subgroup of this compact open subgroup, such that it is neat, and such that it acts trivially on $\mathbb{V}/3\mathbb{V}$. Corresponding to $V$ there is a family of $\mathbb{Q}_{\ell}$ local systems ${}_{{\mathrm{et}}}V_{\ell}$ with $\mathbb{Z}_{\ell}$ sub-systems ${}_{{\mathrm{et}}}\mathbb{V}_{\ell}$. Moreover, there is a filtered flat vector bundle ${}_{{}_{\mathrm{dR}}}V$ over $S_{K}(G,X)$ – indeed we have that ${}_{{}_{\mathrm{dR}}}V$ is defined by ×$\mathcal{E}\times_{G}V$. The filtration is induced by the fact that ${}_{{}_{\mathrm{dR}}}V$ is also canonically isomorphic to $\mathcal{P}\times_{G_{E}}V_{E}$. Now, let $p$ be a prime and let $v\mid p$ be a prime of $E$. By [DLLZ], there is a canonical isomorphism between $(_{{}_{\mathrm{dR}}}V)_{E,v}$ and ${}_{p-_{\mathrm{dR}}}V$, the filtered flat bundle obtained by applying the $p$-adic Riemann-Hilbert correspondence to ${}_{{\mathrm{et}}}V_{p}|_{S_{K}(G,X)_{E,v}}$.
@@
  • ${}_{{\mathrm{et}}}\mathbb{V}_{p}$ is crystalline on $\mathscr{S}_{K}(G,X)_{E_{v}}$ for every finite place $v$ of $\mathcal{O}_{E}[1/N]$.
  • For each finite place $v$ of $\mathcal{O}_{E}[1/N]$, let ${}_{{}_{\mathrm{cris}}}\mathbb{V}/\mathscr{S}_{K}(G,X)_{\mathcal{O}_{E_{v}}}$ denote the Fontaine-Laffaille module corresponding to ${}_{{\mathrm{et}}}\mathbb{V}_{p}$ under the Fontaine-Laffaille correspondence. Then the Kodaira–Spencer map of the Higgs bundle $\operatorname{gr}_{F}(_{{}_{\mathrm{dR}}}\mathbb{V})$ is immersive.
@@
The third condition is proven by Esnault-Groechenig [EG] for large enough primes, and this in particular gives an integral structure to the filtered flat bundle ${}_{p-_{\mathrm{dR}}}V$. Under the canonical isomorphism ${}_{p-_{\mathrm{dR}}}V\rightarrow_{{}_{\mathrm{dR}}}V|_{S_{K}(G,X)_{E_{v}}}$, the integral structures induced by ${}_{{}_{\mathrm{cris}}}\mathbb{V}$ and ${}_{{}_{\mathrm{dR}}}\mathbb{V}$ are the same by [PST].
Finally, using this isomorphism, the final condition follows because the same is true for the Higgs bundle of ${}_{{}_{\mathrm{dR}}}V$ over the generic point. ∎
It follows from Theorem 4.1 that $\mathscr{S}_{K}(G,X)_{\mathcal{O}_{E_{v}}}$ is an integral canonical model. It remains to argue that $\mathscr{S}_{K}(G,X)_{\mathcal{O}_{E_{v}}}$ is the unique model admitting a log smooth compactification. Let $\mathcal{S}$ be another such model. The extension property yields that the identity map over the generic fiber extends to $f:\mathcal{S}\to\mathscr{S}_{K}(G,X)_{\mathcal{O}_{E_{v}}}$, and by purity of the branch locus this map is an open immersion. The inverse image of the boundary of $\bar{\mathcal{S}}$ in the closure $\mathcal{W}\subset\bar{\mathcal{S}}\times\bar{\mathscr{S}}_{K}(G,X)_{\mathcal{O}_{E_{v}}}$ of the graph of $f$ must map to the boundary of $\bar{\mathscr{S}}_{K}(G,X)_{\mathcal{O}_{E_{v}}}$, by the smooth extension property. Thus, the extension is an isomorphism. ∎
@@
Suppose $F$ is a number field, $P/F$ a smooth variety, and $f:Z\to P$ a smooth projective family. Let $f:P^{\prime}\to Y$ be the Stein factorization of the (compactification) of the period map associated to $\mathbb{W}_{\mathbb{Z}}:=Rf_{*}\mathbb{Z}_{Z^{\mathrm{an}}}$ as described in 2.7. Then $f$ is defined over $F$. By spreading out over $\mathcal{O}:=\mathcal{O}_{F}[1/n]$ and inverting finitely many primes, we may assume we have smooth $\mathcal{P},\mathcal{Z}/\mathcal{O}$, a log smooth partial compactification $(\mathcal{P}^{\prime},\mathcal{D})/\mathcal{O}$ of $\mathcal{P}$, and $\mathcal{Y}/\mathcal{O}$ normal finite type flat as well as:
@@
    $f:\mathcal{Z}\to\mathcal{P}$ smooth projective
@@
    $\phi:\mathcal{P}^{\prime}\to\mathcal{Y}$ proper with $\mathcal{O}_{\mathcal{Y}}\xrightarrow{\cong}\mathcal{O}_{\mathcal{P}^{\prime}}$.
@@
    ${\mathcal{T}^{j}}$${\mathcal{P}^{\prime}}$${\mathcal{S}^{j}}$${\mathcal{Y}^{\prime j}}$${\mathcal{Y}}$$\scriptstyle{q^{j}}$$\scriptstyle{t^{j}}$$\scriptstyle{\bar{\phi}}$$\scriptstyle{\pi^{j}}$
@@
    There is a log smooth curve (possibly not irreducible) $(\mathcal{C}^{\prime},\Delta)/\mathcal{O}$ and a morphism $g^{\prime}:(\mathcal{C}^{\prime},\Delta)\to(\mathcal{P}^{\prime},\mathcal{D})$ with $\Delta=(\mathcal{C}^{\prime}\times_{\mathcal{P}^{\prime}}\mathcal{D})^{\mathrm{red}}$ such that $g^{\prime}(\Delta)$ meets every component of $\mathcal{D}$ and the base-change $\mathcal{Z}_{\mathcal{C}}$ (where $\mathcal{C}=\mathcal{C}^{\prime}\setminus\mathcal{E}$) has a log smooth compactification $(\mathcal{Z}^{\prime}_{\mathcal{C}},\mathcal{E})/\mathcal{O}$ and a morphism $f^{\prime}:\mathcal{Z}^{\prime}_{\mathcal{C}}\to\mathcal{C}^{\prime}$ extending the natural one.
@@
    The filtered flat vector bundle $(V,\nabla,F^{\bullet}V)$ on $Y$ extends to $(\mathcal{V},\nabla,F^{\bullet}\mathcal{V})$ on $\mathcal{Y}$, and its pullback to $\mathcal{P}$ is identified with the filtered flat vector bundle underlying $H^{k}_{DR}(\mathcal{Z}/\mathcal{P})$. Moreover, the reduction $(\mathcal{V},\nabla,F^{\bullet}\mathcal{V})_{v}$ has quasifinite Kodaira–Spencer map for every finite place $v$.
@@
This yields part 1(a), and both 2 and 3 follow from 1(b) by Corollary 2.18 and Proposition 2.15 (1(c) is clear from construction). Thus it remains to show the extension property in 1(b), and since by Lemma 2.24 we have the extension in characterstic 0, it suffices to show $\mathcal{Y}_{v}$ is an integral canonical model for each finite place $v$. By Theorem 4.3 we must show for each finite place $v$ of $\mathcal{O}$ that there is a Fontaine–Laffaille module $\mathbf{M}$ on $\mathcal{P}^{\prime}_{v}$ with the required properties. The filtered flat vector bundle underlying the $p$-torsion module of $H^{k}_{DR}(\mathcal{Z}_{v}/\mathcal{P}_{v})$ extends to $\mathcal{P}^{\prime}_{v}$, and so by Proposition 4.9 it extends as a Fontaine–Laffaille module whose underlying filtered flat vector bundle is the pullback of $(\mathcal{V},\nabla,F^{\bullet}\mathcal{V})_{\mathcal{O}_{v}}$ on $\mathcal{Y}_{v}$.∎
6 Application to unramified Borel-extension
In this section, we will apply the extension property for Shimura varieties and images of geometric period maps. To that end, let $F$ be a number field, and let $(Y,\overline{Y},\mathbb{V})$ denote one of the following:
  1. 1.
    $Y\subset\overline{Y}$ is a Shimura variety embedded inside a log smooth toroidal compactification, and $\mathbb{V}$ is ${}_{{\mathrm{et}}}\mathbb{V}_{\ell}$ associated to the adjoint representation of the underlying group $G$ as in Section 5.1. We pick $\ell=3$.
  2. 2.
    $Y\subset\overline{Y}$ is as in Section 5.2. Recall that there is a smooth proper family over a smooth variety $P$, and $Y$ is the image of the period morphism from $P$. For every prime $\ell$, there is an etale $\ell$-adic local system on $P$ induced by the smooth proper family, and this $\ell$-adic local system descends to an $\ell$-adic local system $\mathbb{V}$ on $Y$. As in the Shimura case, we will choose $\ell=3$.
In the non-Shimura setting, we assume the existence of stratified resolution of $\mathcal{Y}/\mathcal{O}_{F}[1/N]$ as in Definition 4.2 where the resolution is also defined over $\mathcal{O}_{F}[1/N]$ where $N$ is some large enough integer. We suppose that $p\nmid N$ is a large enough prime. In particular, we suppose that $p$ is large enough such that the main theorems of Section 5 hold, and such that $p\nmid\#\operatorname{GL}_{n}(\mathbb{Z}/3\mathbb{Z})$, where $n$ is the rank of $\mathbb{V}$.
Definition 6.1.
Let $K$ denote a finite extension of $W(\overline{\mathbb{F}}_{p})[1/p]$ or $\mathbb{F}_{p}((t))$. We say that a point $x\in Y(K)$ has good reduction if its specialization lies in the interior $\mathcal{Y}(\overline{\mathbb{F}}_{p})$. We say that $x$ has bad reduction otherwise.
The content of the following lemma is that a point having good reduction or bad reduction can be tested by considering the fiber of $\mathbb{V}$ at that point.
Lemma 6.2.
Let $K$ be as in Definition 6.1. Let $x\in Y(K)$. Then $x$ has bad reduction if and only if the inertial action of $I_{x}\subset\operatorname{Gal}(\overline{K}/K)$ on $\mathbb{V}_{x}$ is quasi-unipotent of infinite order.
Proof.
By arguments similar to those in [PST] (which we may use, as $\mathbb{V}$ has prime-to-$p$ monodromy given our choice of $p$), we have that a point $x\in Y(K)$ has bad reduction if and only if the inertial action on the fiber at $x$ is quasi-unipotent of infinite order. Indeed, in the Shimura case, this is just [PST, Theorem 4.4] (the result in loc. cit. is for finite extensions of $W(\overline{\operatorname{\mathbb{F}}}_{p})$, but the identical argument goes through in the case of finite extensions of $\overline{\operatorname{\mathbb{F}}}_{p}((t))$). In the geometric case, at the cost of replacing $K$ by a finite extension, there is a point $y\in P(K)$ lifting $x$. The same argument as in [PST, Theorem 4.4] yields that $y$ specializes to the boundary of $P$ if and only if the inertial action is trivial. The claim follows for $x$ as $y$ specializes to the boundary of $Y$ if and only $x$ specializes to the boundary of $P$. ∎
Let $\mathsf{D}$ denote the rigid-analytic closed unit disc (defined over $\mathbb{Q}_{p}$) and let $\mathsf{D}^{\times}=\mathsf{D}\setminus\{0\}$. Recall that a thin annulus is a rigid analytic annulus whose outer and inner radii are the same, and a thick annulus is one whose inner and outer radii are different. We are now ready to state the main theorem of this section – this is a one-dimensional analogue of Borel’s extension theorem in the unramified setting (see the main result of [padicBorel]).
Theorem 6.3.
Let $Y$ be as above, and let $f:\mathsf{D}^{\times}\rightarrow Y^{{\mathrm{an}}}$ denote a rigid-analytic map defined over $W[1/p]$, where $W=W(\overline{\mathbb{F}}_{p})$. Suppose that the image of $f$ contains a point with good reduction. Then, $f$ extends to a rigid-analytic map $\mathsf{D}\rightarrow Y^{{\mathrm{an}}}$.
We first need the following lemma:
Lemma 6.4.
Let $f:\mathsf{D}^{\times}\rightarrow Y^{{\mathrm{an}}}$ be a rigid-analytic map defined over a finite extension of $W[1/p]$, such that the image of $f$ contains a point with good reduction. Then the image of $f$ is entirely contained in the good reduction locus.
Proof.
Note that the local system $f^{-1}\mathbb{V}$ has prime-to-$p$ monodromy. The proof of this lemma is essentially the same as [padicBorel, Theorem 3.3]. As in loc. cit., it suffices to prove the lemma for maps from closed annuli. Proposition 3.6 of loc. cit. proves that an $\ell$-adic local system on a thin annulus must be unramified everywhere if it is at a single point.
To deduce the case of thick annuli from thin annuli, we proceed as in [padicBorel, Section 3.1.2]. The argument applies verbatim to reduce to proving that any map $f:\mathbb{G}_{m}\rightarrow\bar{\mathcal{Y}}_{\mathbb{F}_{p}}$ extends to a map from $\mathbb{P}^{1}$. Note that the local system $\mathbb{V}/\mathcal{Y}_{\mathbb{F}_{p}}$ is point-wise pure. Indeed, in the Shimura case, $\mathbb{V}$ is a direct sum of irreducible local systems with trivial determinant (this follows from the fact that the local system is induced by the adjoint representation of $G$). In the case of geometric period images, it suffices to prove point-wise-purity with $P$ in place of $Y$. But this is true as $\mathbb{V}$ is induced by a geometric family over $P$.
Suppose now that $x\in\mathbb{P}^{1}\setminus\mathbb{G}_{m}$ maps to the boundary. We have that $f^{-1}(\mathbb{V})$ has quasi-unipotent local monodromy of infinite order. We have that $f$ descends to some finite field. In both the cases, the local system originally arises from an irreducible local system over a smooth base, and is therefore point-wise pure over that pure base. As this is a point-by-point condition, it is also point-wise pure over $Y$, and therefore over when pulled back to $\mathbb{G}_{m}$ by $f$. By [DeligneWeil2, Corollaire 3.4.12], the local system has semi-simple geometric monodromy over $\mathbb{G}_{m}$. As the prime-to-$p$ geometric fundamental group of $\mathbb{G}_{m}$ is abelian, the local system pulled back to $\mathbb{G}_{m}$ must have toric geometric monodromy. Therefore, the local monodromy around the boundary of $\mathbb{G}_{m}$ is not quasi-unipotent with infinite order in the first place, and the result follows. ∎
We are now ready to prove Theorem 6.3.
Proof.
By the lemma above, we may assume that the image of $\mathsf{D}^{\times}$ is contained in the good reduction locus of $Y^{{\mathrm{an}}}$. Let $Y^{j}\subset Y$ denote the smallest stratum that contains the image of $\mathsf{D}^{\times}$. We claim that the map $f:\mathsf{D}^{\times}\rightarrow{Y^{j}}^{{\mathrm{an}}}$ lifts to a map $g:\mathsf{D}^{\times}\rightarrow{S^{j}}^{{\mathrm{an}}}$. Indeed, $\mathsf{D}^{\times}\cap{Y^{j-1}}^{{\mathrm{an}}}$ is a analytic closed subset, and so consists of a sequence of points $Q_{i}\in\mathsf{D}^{\times}$ with the property that only finitely many of the $Q_{i}$ lie in any closed sub-annulus of $\mathsf{D}^{\times}$. Pick $D_{i}$ to be a small closed disc centered around $Q_{i}$ such that $D_{i}\cap Y^{j-1}=Q_{i}$. It suffices to show that the map $D_{i}\rightarrow{Y^{j}}^{{\mathrm{an}}}$ lifts to $D_{i}\rightarrow S^{j}$. But this follows from the fact that $S^{j}\rightarrow Y^{j}$ is proper.
Let $\mathsf{A}_{n}$ denote the rigid-analytic annulus with outer radius 1 and inner radius $\frac{1}{p^{n}}$, and let $g_{n}=g|_{\mathsf{A}_{n}}$. Therefore, the map $f_{n}$ is induced by a map of formal schemes $\mathscr{X}_{n}\rightarrow\widehat{\mathcal{S}^{j}}^{p}$, where $\mathscr{X}_{n}$ is a formal scheme with rigid generic fiber $\mathsf{A}_{n}$, and $\widehat{\mathcal{S}^{j}}^{p}$ is the $p$-adic completion of $\mathcal{S}^{j}$. By [padicBorel, Appendix], we may assume that the special fiber of $\mathscr{X}_{n}$ is a union of projective lines and two affine lines, and that the the irreducible components form a tree. As $\mathscr{X}_{n}\mod p^{2}$ is a scheme that maps algebraically to $\mathscr{S}_{K}(G,X)$ over $W/p^{2}$, the Fontaine-Laffaille mod module mod $p$ pulls back to the special fiber of $\mathscr{X}_{n}$. In the Shimura case, Corollary 3.14 yields that every every $\mathbb{P}^{1}$ contracts to a point. In the case of the image of period maps, an argument similar to the claim in the proof of Proposition 4.4 applies to show that the map $\pi^{j}\circ g$ contracts every $\mathbb{P}^{1}$.
Therefore we may assume that the special fiber of $\mathscr{X}_{n}$ is just $\overline{\mathbb{F}}_{p}[x,y]/(xy)$. Further, we have that the tube over the node is the open annulus $\mathsf{A}_{n}^{\circ}\subset\mathsf{A}_{n}$. Therefore, we have that $f(\mathsf{A}_{n}^{\circ})$ is contained in the residue ball over some fixed $z\in\mathcal{Y}^{j}(\overline{F}_{p})$. It follows that the image of the open punctured unit disk $\mathsf{D}^{\circ,\times}\subset\mathsf{D}^{\times}$ is contained in the residue ball over $z\in\mathcal{Y}^{j}(\overline{\mathbb{F}}_{p})$, and thus is bounded. We now apply the rigid-analytic Riemann extension theorem to conclude that $f$ extends to a map $\mathsf{D}\rightarrow Y^{{\mathrm{an}}}$.

### v2

@@ 4.3 Making period maps proper
Let $F/\mathbb{Q}_{p}$ be a finite unramified extension with ring of integers $\mathcal{O}$. Let $\mathcal{P}/\mathcal{O}$ be smooth with a partial log smooth compactification $(\mathcal{P}^{\prime},\mathcal{D})/\mathcal{O}$. Let $f:\mathcal{Z}\to\mathcal{P}$ be a smooth projective morphism with geometrically connected fibers of relative dimension $<p$. Let $\mathbf{M}=H^{k}_{DR}(\mathcal{Z}/\mathcal{P})$ be the Fontaine–Laffaille module with underlying filtered flat vector bundle the De Rham cohomology of $\mathcal{Z}/\mathcal{P}$. Assume:
@@
    There is a generically finite proper morphism $g:(\mathcal{Q},\mathcal{E})\to(\mathcal{P}^{\prime},\mathcal{D})$ from a log smooth $(\mathcal{Q},\Delta)/\mathcal{O}$ and a semistable morphism $(\mathcal{W},\Delta)\to(\mathcal{Q},\mathcal{E})$ from a log smooth $(\mathcal{W},\Delta)/\mathcal{O}$ extending the base-change of $f$ (possibly after shrinking $\mathcal{P}$).
@@
    $g^{\prime*}(M^{\prime},\nabla^{\prime},F^{\bullet}M^{\prime})\cong H^{k}_{DR}((\mathcal{W},\Delta)/(\mathcal{P}^{\prime},\mathcal{D}))$ via an isomorphism extending the natural one, where $H^{k}_{DR}((\mathcal{W},\Delta)/(\mathcal{P}^{\prime},\mathcal{D}))$ is the relative log De Rham cohomology.
@@
According to [faltings, Thm 6.2], the completion of the filtered flat bundle $H^{k}_{DR}((\mathcal{W},\Delta)/(\mathcal{P}^{\prime},\mathcal{D}))$ naturally underlies a log Fontaine–Laffaile module [faltings, §2i]. By assumption the connection has trivial residue, so it follows it is in fact a Fontaine–Laffaille module. Now apply Lemma 4.8(2). ∎
@@ 5 Existence of integral canonical models
Let $V$ denote a faithful rational representation of $G$, and $\mathbb{V}\subset V$ denote a lattice. The choice of $\mathbb{V}$ specifies a compact open subgroup of $G(\mathbb{A}_{f})$ by considering the stabilizers of $\mathbb{V}_{\hat{\mathbb{Z}}}$. We choose the level structure $K$ to be a subgroup of this compact open subgroup, such that it is neat, and such that it acts trivially on $\mathbb{V}/3\mathbb{V}$. Corresponding to $V$ there is a family of $\mathbb{Q}_{\ell}$ local systems ${}_{{\mathrm{et}}}V_{\ell}$ with $\mathbb{Z}_{\ell}$ sub-systems ${}_{{\mathrm{et}}}\mathbb{V}_{\ell}$. Moreover, there is a filtered flat vector bundle ${}_{{}_{\mathrm{dR}}}V$ over $S_{K}(G,X)$ – indeed we have that ${}_{{}_{\mathrm{dR}}}V$ is defined by ×$\mathcal{E}\times_{G}V$. The filtration is induced by the fact that ${}_{{}_{\mathrm{dR}}}V$ is also canonically isomorphic to $\mathcal{P}\times_{G_{E}}V_{E}$. Now, let $p$ be a prime and let $v\mid p$ be a prime of $E$.
@@
  • For each finite place $v$ of $\mathcal{O}_{E}[1/N]$, ${}_{{\mathrm{et}}}\mathbb{V}_{p}(m)$ is crystalline on $\mathscr{S}_{K}(G,X)_{E_{v}}$ for $m$ sufficiently small so that $V(m)$ has positive Hodge degrees.
  • For each finite place $v$ of $\mathcal{O}_{E}[1/N]$, let ${}_{{}_{\mathrm{cris}}}\mathbb{V}/\mathscr{S}_{K}(G,X)_{\mathcal{O}_{E_{v}}}$ denote the Fontaine-Laffaille module corresponding to ${}_{{\mathrm{et}}}\mathbb{V}_{p}$ under the Fontaine-Laffaille correspondence. Then the Griffiths bundle of ${}_{{}_{\mathrm{cris}}}\mathbb{V}$ is ample.
@@
The third condition is [PST, Thm 7.1]
The final condition follows because the same is true for the filtered flat bundle ${}_{{}_{\mathrm{dR}}}V$ over the generic point. ∎
It follows from Theorem 4.1 that $\mathscr{S}_{K}(G,X)_{\mathcal{O}_{E_{v}}}$ is an integral canonical model. It remains to argue that $\mathscr{S}_{K}(G,X)_{\mathcal{O}_{E_{v}}}$ is the unique model admitting a log smooth compactification. Let $\mathcal{S}$ be another such model. The extension property yields that the identity map over the generic fiber extends to $f:\mathcal{S}\to\mathscr{S}_{K}(G,X)_{\mathcal{O}_{E_{v}}}$, and by purity of the branch locus this map is an open immersion. Let $\mathcal{W}\subset\bar{\mathcal{S}}\times\bar{\mathscr{S}}_{K}(G,X)_{\mathcal{O}_{E_{v}}}$ be the closure of the graph of $f$. Over the generic point, the inverse image of the boundary of $\bar{S}$ in $W$ must map to the boundary of $\bar{\mathscr{S}}_{K}(G,X)_{\mathcal{O}_{E_{v}}}$, since every boundary divisor has infinite local monodromy, and the same is therefore true for $\mathcal{W}$. Thus, the extension is an isomorphism. ∎
@@
Suppose $F$ is a number field, $P/F$ a smooth variety, and $f:Z\to P$ a smooth projective family. Let $h_{F}:P^{\prime}\to Y$ be the Stein factorization of the (compactification) of the period map associated to $\mathbb{W}_{\mathbb{Z}}:=Rf_{*}\mathbb{Z}_{Z^{\mathrm{an}}}$ as described in 2.7. Then $h_{F}$ is defined over $F$. By spreading out over $\mathcal{O}:=\mathcal{O}_{F}[1/n]$ and inverting finitely many primes, we may assume we have smooth $\mathcal{P},\mathcal{Z}/\mathcal{O}$, a log smooth partial compactification $(\mathcal{P}^{\prime},\mathcal{D})/\mathcal{O}$ of $\mathcal{P}$, and $\mathcal{Y}/\mathcal{O}$ normal finite type flat as well as:
@@
    $f:\mathcal{Z}\to\mathcal{P}$ smooth projective.
@@
    $h:\mathcal{P}^{\prime}\to\mathcal{Y}$ proper with $\mathcal{O}_{\mathcal{Y}}\xrightarrow{\cong}h_{*}\mathcal{O}_{\mathcal{P}^{\prime}}$.
@@
    ${\mathcal{T}^{j}}$${\mathcal{P}^{\prime}}$${\mathcal{S}^{j}}$${\mathcal{Y}^{\prime j}}$${\mathcal{Y}}$$\scriptstyle{q^{j}}$$\scriptstyle{t^{j}}$$\scriptstyle{h}$$\scriptstyle{\pi^{j}}$
@@
    By [semistab_reduction] there is a semistable model $(\mathcal{W},\Delta)\to(\mathcal{Q},\mathcal{E})$ for the family $f$ up to a smooth alteration $g:(\mathcal{Q},\mathcal{E})\to(\mathcal{P}^{\prime},\mathcal{D})$.
@@
    The filtered flat vector bundle $(V,\nabla,F^{\bullet}V)$ on $Y$ extends to $(\mathcal{V},\nabla,F^{\bullet}\mathcal{V})$ on $\mathcal{Y}$, and its pullback to $\mathcal{P}$ is identified with the filtered flat vector bundle underlying $H^{k}_{DR}(\mathcal{Z}/\mathcal{P})$. Moreover, $(\mathcal{V},\nabla,F^{\bullet}\mathcal{V})$ has ample Griffiths bundle by Theorem 2.28.
@@
This yields part 1a, and both 2 and 3 follow from 1b by Corollary 2.19 and Proposition 2.16. Part 4 is clear from construction. Thus it remains to show the extension property in 1b, and since by Lemma 2.25 we have the extension in characterstic 0, it suffices to show $\mathcal{Y}_{v}$ is an integral canonical model for each finite place $v$. By Theorem 4.3 we must show for each finite place $v$ of $\mathcal{O}$ that there is a Fontaine–Laffaille module $\mathbf{M}$ on $\mathcal{P}^{\prime}_{v}$ with the required properties. The filtered flat vector bundle underlying $H^{k}_{DR}(\mathcal{Z}_{v}/\mathcal{P}_{v})$ extends to $\mathcal{P}^{\prime}_{v}$, and so by Proposition 4.9 it extends as a Fontaine–Laffaille module whose underlying filtered flat vector bundle is the pullback of $(\mathcal{V},\nabla,F^{\bullet}\mathcal{V})_{\mathcal{O}_{v}}$ on $\mathcal{Y}_{v}$.∎